OpenStudy (nerdgirl):

1. You place 3.52 g of sodium hydroxide to dissolve in a calorimeter cup. The cup contains 1,000.0 mL of water at 20.00°C (in a calorimeter cup). As the sodium hydroxide dissolves, the temperature of the surrounding water increases to 23.70°C. Complete the table below and determine the total energy released. Use the formula .

OpenStudy (nerdgirl):

Use the formula Q = m x c x T

OpenStudy (nerdgirl):

HELP!!! I just need help with the first question!

OpenStudy (nerdgirl):

@runninthruthe6 @Smita12 @maddiegirl @Ziferot @Lilium @kidrah69 @nich01a5

OpenStudy (anonymous):

When we are looking for q(soln) we use the TOTAL mass of both. ALSO, we use water's specific heat 1000ml⋅1g/mL=1000g ----> 1000g+3.25g=1003.25g(total) Plug it in: q(soln)=1003.25g⋅4.184J/g°C⋅(23.7°C-20°C...‡ = 15531.1126J released 15531.1126J/3.52g=4412.248J(released)/g Next, enthalpy is q(rxn)/g(usually mol), so we must use q(soln) = -q(rxn) ∴ ΔH=-4412.248J/g

OpenStudy (anonymous):

lol answers.com

OpenStudy (anonymous):

hen we are looking for q(soln) we use the TOTAL mass of both. ALSO, we use water's specific heat 1000ml⋅1g/mL=1000g ----> 1000g+3.25g=1003.25g(total) Plug it in: q(soln)=1003.25g⋅4.184J/g°C⋅(23.7°C-20°C...‡ = 15531.1126J released 15531.1126J/3.52g=4412.248J(released)/g Next, enthalpy is q(rxn)/g(usually mol), so we must use q(soln) = -q(rxn) ∴ ΔH=-4412.248J/g

OpenStudy (nerdgirl):

Thanks to you both. :) So does -4,412.248J/g go in the two empty spaces in Question 1?

OpenStudy (anonymous):

yea

OpenStudy (nerdgirl):

Cool thanks. :) How about the table?

OpenStudy (anonymous):

complete different story im lost lol

OpenStudy (nerdgirl):

I'm talking about the table in question 1 - how do I get the values to plug in there?

OpenStudy (anonymous):

srry im rally tired zzzzzzzzzzzzzzzzzzzzzzzzzzzz